OpenStudy (anonymous):

Can any kind soul please help me with this problem? What is the volume of the rectangular prism? A. 896 ft3 B. 30 ft3 C. 224 ft3 D. 576 ft3

jimthompson5910 (jim_thompson5910):

Volume of Prism = (Area of Base) * (Height)

OpenStudy (anonymous):



OpenStudy (anonymous):

14*8*8/6 ?

jimthompson5910 (jim_thompson5910):

Area of Base = ??

OpenStudy (anonymous):

/3

OpenStudy (anonymous):

wait no that's not right

OpenStudy (anonymous):

Volume =H*L*W =14*8*8 =64*14=896feet^3 Answer:(a)

OpenStudy (anonymous):

Thank you:)